# Can light-rings self-gravitate?

### v1

@@ Abstract
I Introduction
Light-rings, namely, the location of unstable circular orbits, play a crucial role in the study of astrophysical compact objects as shown by a number of a theoretical investigations exploring the spacetime on those scales (see, e.g., Chirenti and Rezzolla (2007); Vincent et al. (2016); Glampedakis et al. (2017); Cardoso et al. (2017); Jai-akson et al. (2017); Glampedakis and Pappas (2018); Cardoso et al. (2016); Carballo-Rubio et al. (2018); Cardoso and Pani (2019); Völkel et al. (2021); Kocherlakota et al. (2021); Völkel et al. (2022); Eichhorn et al. (2023); Moreira et al. (2024); Pedrotti and Vagnozzi (2024); Genzel et al. (2024)), but also by the observations of supermassive black holes Akiyama et al. (2019, 2022a, 2022b); Kocherlakota and Rezzolla (2022) or through the ringdown associated with the gravitational-wave signal from merging black-hole binaries Abbott and et al (2016); Abbott et al. (2019, 2021a, 2021b). Indeed, it is generally accepted that the study of light-rings, either via black-hole imaging or via black-hole quasi-normal modes, represents one of the most promising routes to test theories of gravity beyond general relativity.
Interestingly, while light-rings in general relativity are well understood, both for nonrotating and rotating black holes, but also for other compact objects not possessing an event horizon, and have been the subject of uncountable works, there is an question that so far has escaped the attention of the literature, namely, “can light-rings self-gravitate?”
@@ I Introduction
However, as the number of photons in the light-ring increases with time, its mass will no longer be negligible and hence it will start to “self-gravitate”, that is, it will have a sizable impact on the background spacetime. Under these conditions, the dynamics of the photons can no longer be considered as that in the original black-hole spacetime in vacuum, but it will be regulated by the geometry of the spacetime resulting from the proper solution of the Einstein equations for a black hole surrounded by the energy contributions coming from the thin null shells of photons on (unstable) circular orbits.
@@
The structure of the paper is as follows. Section II briefly reviews the properties of light-rings around Schwarzschild black holes when these are seen as populated by test photons and illustrates the solution of the Einstein equations to construct a hierarchy of discrete, thin, self-gravitating light-rings, pointing out its intrinsic instability. Similarly, Sec. III discusses how to solve the Einstein equations to construct a continuous self-gravitating light-region and it shows that it is equally unstable. Finally, Sec. IV summarises our results and the implications for realistic black holes.
II Discrete sequence of light-rings
We start by recalling that for an axisymmetric and stationary spacetime described by spherical polar coordinates $(t,r,\theta,\phi)$, $\partial_{t}$ and $\partial_{\phi}$ are Killing vectors. Light-rings in such axisymmetric and stationary spacetimes containing a black hole are then defined as regions of the spacetime in which null geodesics only have momenta along such Killing vectors, namely, along the $t$ and $\phi$ directions. If we further restrict to a spherically symmetric configuration, which is the interest of this paper, we can write the metric as
@@ II Discrete sequence of light-rings
and the locations of the light-rings can be obtained by studying the Hamiltonian for a massless particle with four-momentum $p_{\mu}$
$\mathcal{H}:=g^{\mu\nu}p_{\mu}p_{\nu}=g^{rr}p_{r}^{2}+g^{\theta\theta}p_{\theta}^{2}+V(r,\theta)=0\,.$ (2)
More specifically, given the potential $V(r,\theta)$
$V(r,\theta):=\frac{E^{2}}{g_{tt}}+\frac{\Phi^{2}}{r^{2}\sin^{2}\theta}\,,$ (3)
where $E:=-p_{t}$ and $\Phi:=p_{\phi}$ are constants of motion, the definition (2) and Hamilton’s equations
@@
where the dot refers to a derivative with respect to the affine parameter, imply that the conditions for the occurrence of light-rings, i.e., $p_{r}=p_{\theta}=\dot{p}_{\mu}=0$, are equivalent to $V(r,\theta)=0$ and $\partial_{\mu}V(r,\theta)=0$. Furthermore, motion on the light-ring will be stable if its location corresponds to either a minimum of the potential and unstable if it corresponds to a maximum or a saddle point.
Rather than considering the potential, we can determine the position of the light-rings by restricting to the equatorial plane and studying the stationary points of the effective potential (see e.g., Cunha et al. (2016, 2017); Di Filippo (2024)) $H(r):={g_{tt}}/{g_{\phi\phi}}$. This approarch reveals, for example, that a single light-ring exists in the case of a Schwarzschild black hole and it is located at $r=3M$, thus marking the position of the only circular but unstable photon orbit.
Our next step consists of adding a discrete set of shells on a Schwarzschild background. In particular, in our “Gedankenexperiment” we imagine that the advanced civilisation mentioned above is able to populate the light-ring of a Schwarzschild black hole of initial mass $M_{0}$ with photons having a fine-tuned impact parameter that allows their orbits to be circular at the light-ring. We assume that the angular components of the momentum of each photon are randomly but uniformly distributed, so that the corresponding light-ring will be be a 2-sphere, thus preserving the spherical symmetry of the spacetime. As the number of photons in the light-ring increases, the energy of the shell of photons will no longer be negligible, impacting the background spacetime and hence leading to a new spacetime metric that is no longer that of the Schwarzschild black hole with mass $M_{0}$. In order to obtain the resulting metric, we assume that the new metric is actually composed of a black hole and a discrete series of $N$ nested null shells of photons on (unstable) circular orbits. This allows us to consider a fully generic approach but it is trivial to restrict the relevant results to the case of a single shell, i.e., for $N=1$.
@@
where $n\in\{0,N\}$ and $\xi^{2}_{n}$ is a set of coefficients needed to guarantee the continuity between the different portions of the spacetime and where $\xi^{2}_{0}=1$. Clearly, the region $n=0$ corresponds to the section between the black hole and the first shell, while the $n$-th region corresponds to the section between the $(n-1)$-th and the $n$-th shell. Obviously, the region outside the $n$-th shell is described by a Schwarzschild spacetime with mass $M_{n}$.
The mass parameter $M_{n}$ in Eq. (5) represents the total mass enclosed within the corresponding region and is thus given by
$M_{n}:=M_{0}+\sum_{i=1}^{n}\delta m_{i}=M_{0}+\delta M_{n}\,,$ (6)
where $\delta m_{i}$ indicates the mass of a single shell and $\delta M_{n}$ denotes the sum of the masses of the first $n$ shells, i.e.,
$\delta M_{n}:=\sum_{i=1}^{n}\delta m_{i}\,.$ (7)
Figure 1: Shown with a black solid line is the effective potential in the case of a discrete hierarchy of light-rings (left panel, where $M_{0}=1$, $\delta m_{1}=\delta m_{2}=0.01$ and the light rings are at $r_{0}=3\,M_{0}$, $r_{1}=3\left(M_{0}+\delta m_{1}\right)$ and $r_{2}=3\left(M_{0}+\delta m_{1}+\delta m_{2}\right)$) or a continuous sequence of light-rings (right panel, where $M_{0}=1$, $R_{\rm in}=3.0\,M_{0}$ and $R_{\rm out}=3.1\,M_{0}$). The location of the light rings is shown with vertical dashed lines in the left panel and with a shaded area in the right panel. In both cases, the red dashed line shows the effective potential for a Schwarzschild black hole.
As mentioned above, $\xi^{2}_{n}$ are constants and can be fixed by simply redefining the time coordinate in each spacetime region. While this is in principle possible, it would complicate the subsequent analysis and hence we determine these constants by requiring the requiring the $\mathcal{C}_{0}$ continuity of the induced metric on the shells.
@@
whose extremal points marking the position of the unstable photon circular orbits can be easily computed after imposing $\partial_{r}H_{n}(r)=0$ and are located at the $n$ different shells with coordinate radii $r_{n}=3M_{n}$. Note that the location of the $n$-th shell coincides with the location of the light-ring of the $n-1$ spacetime region, i.e., $R_{n}=r_{n-1}$, and this allows us to determine the conditions under which the hierarchy of light-rings actually forms a new black hole. In practice, this would happen for $M_{n}/R_{n}\geq 1/2$, that is, for
$\frac{M_{n}}{R_{n}}=\frac{M_{n}}{3M_{n-1}}=\frac{M_{n-1}+\delta m_{n}}{3M_{n-1}}\geq\frac{1}{2}\,.$ (10)
Stated differently, no new black hole is formed as long as the additional mass outside the initial black hole is
$\frac{\delta m_{n}}{M_{0}}\leq\frac{M_{n-1}}{2M_{0}}=\frac{1}{2}\left(1+\frac{\delta M_{n-1}}{M_{0}}\right)\,,$ (11)
which is easily true since we are exploring a scenario with $\delta m_{n}/M_{0}\ll 1$.
@@
$\xi^{2}_{1}=\frac{R_{1}-{2M_{0}}}{R_{1}-{2M_{1}}}\,,$ (12)
$r_{1}=3M_{1}=3\left(M_{0}+\delta M_{1}\right)\,,$ (13)
@@
Is this configuration stable? In principle, to address this question we would need to perform a linear perturbation analysis taking Eq. (5) as background spacetime and assess the properties of the imaginary eigenfrequencies (see, e.g., Nagar and Rezzolla (2005)). In practice, we take a shortcut and note that it is not difficult to realise that such a spacetime is dynamically unstable. Consider, for instance, the addition of a single photon on any of the shells of the hierarchy. This will be sufficient to change the mass of the corresponding shell and hence place on unstable photon orbits all the photons on the light-rings that are at larger radii than the one on which the photon is deposited. Similarly, if a single photon is absorbed by the black hole, this would destabilise all of the light-rings, dispersing or absorbing all the photons that have been accumulated. learly, while we believe this conclusion is correct, a proper perturbative analysis is needed to confirm it.
III Continuous sequence of light-rings
@@
As mentioned in the introduction, the “Gedankenexperiment” can be carried out also in a different way. More specifically, rather than accumulating photons on the first light-ring till they produce a discrete contribution to the spacetime and generate a new light-ring, we can imagine a situation in which the advanced civilization concentrates on adding photons in the right amount, both in terms on numbers and impact parameters, so that the spacetime properties are varied smoothly. This would correspond to a situation in which both the mass of each individual shell and the distance among shells goes to zero while the total mass of the shells remains finite. This configuration would then describe a continuous energy distribution within a finite region of light-rings, i.e., a “light-region” with radius $r\in\left[R_{\rm in},R_{\rm out}\right]$, with $R_{\rm in}=3\,M_{0}$.
Hence, we start by replacing the discrete nature of the mass in the light-ring given by Eq. (6) with the continuous function
$M(r)=M_{0}+\delta M(r)\,,$ (15)
@@
where $\delta M(r)$ denotes the full energy within a radius $r$, so that $\delta M(R_{\rm in})=0$. Furthermore, as $M(r)$ denotes the total energy within a radius $r$, it must be a non-decreasing function, i.e., $\partial_{r}\delta M(r)\geq 0$. Also in this case we can use the coefficients $\xi_{n}$ and calculate their values in the continuous limit; it is however more straightforward to use a perturbative approach and expanding in terms of the ratio $\delta M/M\ll 1$. More specifically, we can rewrite Eq. (8) as
$\begin{array}[]{rl}\xi^{2}_{n}=&\displaystyle\left(\frac{R_{n}-{2M_{n-1}}}{R_{n}-{2M_{n}}}\right)\xi^{2}_{n-1}=\prod_{i=1}^{n}\left(1+\frac{2\delta m_{i}}{R_{i}-{2M_{i}}}\right)\\ =&\displaystyle 1+\sum_{i=1}^{n}\frac{2\delta m_{i}}{R_{i}-{2M_{i}}}+\\ &\displaystyle\frac{1}{2}\sum_{i,j=1}^{n}\left(\frac{2\delta m_{i}}{R_{i}-{2M_{i}}}\right)\left(\frac{2\delta m_{j}}{R_{j}-{2M_{j}}}\right)+\mathcal{O}\left(\left(\frac{\delta M}{M}\right)^{3}\right)\,,\end{array}$
so that
$\xi^{2}_{n}\approx 1+\sum_{i=1}^{n}\frac{2\delta m_{i}}{R_{i}-{2M_{i}}}+\frac{1}{2}\left(\sum_{i=1}^{n}\frac{2\delta m_{i}}{R_{i}-{2M_{i}}}\right)^{2}\,,$ (16)
when neglecting terms $\mathcal{O}\left(\left({\delta M}/{M}\right)^{3}\right)$ and where quadratic term is necessary since the effective potential does not vary at $\mathcal{O}\left(\left({\delta M}/{M}\right)\right)$. The continuous limit of expression (16) can be obtained by replacing the sum with an integral, namely
$\xi^{2}(r)=\displaystyle 1+\int_{3M_{0}}^{r}\frac{2M^{\prime}(\bar{r})}{\bar{r}-2M(\bar{r})}{\rm d}\bar{r}+\frac{1}{2}\left(\int_{3M_{0}}^{r}\frac{2M^{\prime}(\bar{r})}{\bar{r}-2M(\bar{r})}{\rm d}\bar{r}\right)^{2}\,,$ (17)
whose analytic solution requires now a reasonable assumption. In particular, after defining $\delta r:=r-3M_{0}$, we can rewrite the linear term in Eq. (17) as
$\int_{3M_{0}}^{r}\frac{2M^{\prime}(\bar{r})}{\bar{r}-2M(\bar{r})}{\rm d}\bar{r}=$ $\int_{3M_{0}}^{r}\frac{2M^{\prime}(\bar{r})}{3M_{0}+\delta r-2M(\bar{r})}{\rm d}\bar{r}$
$=$ $\int_{M_{0}}^{M_{0}+\delta M}\frac{2{\rm d}M}{3M_{0}+\delta r(\delta M)-2M}\,,$ (18)
where we have indicated with $\delta r(\delta M)$ the inverse of the function $\delta M(r)$. Making the assumption (whose correctness we will validate a-posteriori) that $\delta M$ and $\delta r$ are linearly, related, i.e., that $\delta r(\delta M)=\alpha\delta M$, with $\alpha$ a positive constant, we can solve the two integrals in (17) and obtain
$\xi^{2}(r)=1+\frac{2\delta M(r)}{M_{0}}+\frac{\left(4-\alpha\right)\delta M^{2}(r)}{\left(M_{0}\right)^{2}}+\mathcal{O}\left(\left(\frac{\delta M}{M}\right)^{3}\right)\,.$ (19)
As a result, at the quadratic order in $\delta M/M$ the effective potential reads
$H(r)$ $=\frac{1}{r^{2}}\left(1-\frac{2M_{0}}{r}\right)+\frac{2(r-3M_{0})}{r^{3}}\left(\frac{\delta M}{M_{0}}\right)$
$\phantom{=}+\frac{2M_{0}(\alpha-6)+r(4-\alpha)}{r^{3}}\left(\frac{\delta M}{M_{0}}\right)^{2}\,,$ (20)
where it should be noted that also the second term in (III) is quadratic in $\delta M/M$ being the product of a term $\delta r/M_{0}$ and of a term $\delta M/M$. Hence, as anticipated, the first modification of the effective potential arises at the quadratic order in $\delta M/M$ Requiring now that for the light-ring $H(r)={\rm const.}=H(R_{\rm in})$, yields
$\delta M=\displaystyle\frac{\sqrt{9-3\alpha}+3}{3\alpha}\delta r+\mathcal{O}\left(\delta r^{2}\right)\,.$ (21)
This result is compatible with our linearity assumption $\delta r(\delta M)=\alpha\delta M$ in Eq. (21) provided that the value of the constant is $\alpha=3$. In summary, we have shown that it is possible to create a self-gravitating “light-region” that extends from the position of the original light-ring $R_{\rm in}$ up to $R_{\rm out}$ and whose effective potentials reads
$H(r)=$ $\frac{1}{r^{2}}\left(1-\frac{2M_{0}}{r}\right)+\frac{2(r-3M_{0})}{r^{3}}\left(\frac{\delta M}{M_{0}}\right)$
$-\frac{6M_{0}-r}{r^{3}}\left(\frac{\delta M}{M_{0}}\right)^{2}\,.$ (22)
@@
The corresponding potential is shown with a black solid line in the right panel of Fig. 1 and it is then easy to appreciate how we have essentially created a smooth region of light-rings where the potential is flat at the order considered in the region $R_{\rm in}\leq r\leq R_{\rm out}$, hence leading to an infinite sequence of neutral circular photon orbits in such a region, i.e., with zero first and second derivative of the effective potential.
@@
We can now ask the same question: is this configuration stable? Obviously, a photon on a circular orbit well inside the light-region, i.e., at $R_{\rm in}\lesssim r\lesssim R_{\rm out}$, and that is perturbed, would still encounter in its neighbourhood other circular orbits and hence remain in the newly and perturbed location. However, it is also easy to conclude that the whole light-region is dynamically unstable after a small change in the mass of the black hole. Consider in fact that the mass of the spacetime is changed by a small but finite amount $\delta M_{0}$ without changing the mass in the light-rings, namely, consider the new mass profile $M(r)=M_{0}+\delta M_{0}+\delta M(r)$ with both $M_{0}$ and $\delta M_{0}$ constants. The new effective potential will be
$H(r)$ $=\frac{1}{r^{2}}\left(1-\frac{2M_{0}}{r}\right)+\frac{2(r-3M_{0})}{r^{3}}\left(\frac{\delta M}{M_{0}}\right)$
$\phantom{=}-\frac{6M_{0}-r}{r^{3}}\left(\frac{\delta M}{M_{0}}\right)^{2}-\frac{2\delta M_{0}}{r^{3}}\,,$ (23)
@@
where the first three terms, which are the same appearing in Eq. (III), have a combined zero radial derivative, so that, in practice: $\partial_{r}H(r)={6\delta M_{0}}/{r^{4}}>0$. Stated differently, a small perturbation in the central black hole would completely destabilise the circular orbits in the light-region, dispersing all the photons carefully accumulated there.
IV Conclusion
@@
Light-rings represent one of the most interesting aspects of the spacetime around compact objects in general and of black holes in particular. Being characterised by circular but unstable orbits, they represent a powerful tool to probe the spacetime near these objects and test different theories of gravity, both in terms of electromagnetic and gravitational radiation. So far, however, the self-gravity of light-rings has been neglected and this is because one expects that, on average, the number of photons entering the light-ring equals that of photons leaving the light-ring, and that such a number is not sufficient to modify the spacetime appreciably. We have here explored what would happen when this assumption is not true and for this we considered ways in which the light-ring in a Schwarzschild spacetime can be filled with carefully injected photons so as to make the light-ring self-gravitate.

### v2

@@ Abstract
Introduction.  Light-rings, namely, the location of unstable circular orbits, play a crucial role in the study of astrophysical compact objects as shown by a number of a theoretical investigations exploring the spacetime on those scales (see, e.g., Chirenti and Rezzolla (2007); Vincent et al. (2016); Glampedakis et al. (2017); Cardoso et al. (2017); Jai-akson et al. (2017); Glampedakis and Pappas (2018); Cardoso et al. (2016); Carballo-Rubio et al. (2018); Cardoso and Pani (2019); Völkel et al. (2021); Kocherlakota et al. (2021); Völkel et al. (2022); Eichhorn et al. (2023); Moreira et al. (2024); Pedrotti and Vagnozzi (2024); Genzel et al. (2024)), but also by the observations of supermassive black holes Akiyama et al. (2019, 2022a, 2022b); Kocherlakota and Rezzolla (2022) or through the ringdown associated with the gravitational-wave signal from merging black-hole binaries Abbott and et al (2016); Abbott et al. (2019, 2021a, 2021b). Indeed, it is generally accepted that the study of light-rings, either via black-hole imaging or via black-hole quasi-normal modes, represents one of the most promising routes to test theories of gravity beyond general relativity.
Interestingly, while light-rings in general relativity are well understood, both for nonrotating and rotating black holes Khoo and Ong (2016); Cunha and Herdeiro (2020); Hod (2023), but also for other compact objects not possessing an event horizon Cunha et al. (2017a); Franzin et al. (2024); Di Filippo (2024), and have been the subject of uncountable works, including works aimed at the study of populated spacetimes Andréasson et al. (2017); Andréasson (2021); Macedo et al. (2024), there is an question that so far has escaped the attention of the literature, namely, “can light-rings self-gravitate?”
@@
However, as the number of photons in the light-ring increases with time, and ignoring any photon-photon interaction, its mass will no longer be negligible and hence it will start to “self-gravitate”, that is, it will have a sizable impact on the background spacetime [^1] Photon-photon will tend to destabilise photons previously present on the light-ring, but as long as the lost photons are replaced by new and more numerous photons, the mass of the light-ring will grow.. Under these conditions, the dynamics of the photons can no longer be considered as that in the original black-hole spacetime in vacuum, but it will be regulated by the geometry of the spacetime resulting from the proper solution of the Einstein equations for a black hole surrounded by the energy contributions coming from the thin null shells of photons on (unstable) circular orbits.
@@
Discrete sequence of light-rings.  We start by recalling that for an axisymmetric and stationary spacetime described by spherical polar coordinates $(t,r,\theta,\phi)$, $\partial_{t}$ and $\partial_{\phi}$ are Killing vectors. Light-rings in such axisymmetric and stationary spacetimes containing a black hole are then defined as regions of the spacetime in which null geodesics only have momenta along such Killing vectors, namely, along the $t$ and $\phi$ directions. If we further restrict to a spherically symmetric configuration, which is the interest of this paper, we can write the metric as
@@
and the locations of the light-rings can be obtained by studying the Hamiltonian for a massless particle with four-momentum $p_{\mu}$, $\mathcal{H}:=g^{\mu\nu}p_{\mu}p_{\nu}$. More specifically, given the potential $V(r,\theta)$
$V(r,\theta):=\frac{E^{2}}{g_{tt}}+\frac{L^{2}}{r^{2}\sin^{2}\theta}\,,$ (2)
where $E:=-p_{t}$ and $L:=p_{\phi}$ are constants of motion, the Hamilton equations read
@@
where the dot refers to a derivative with respect to the affine parameter, imply that the conditions for the occurrence of light-rings, i.e., $p_{r}=p_{\theta}=\dot{p}_{\mu}=0$, are equivalent to $V(r,\theta)=0$ and $\partial_{\mu}V(r,\theta)=0$. Furthermore, motion on the light-ring will be stable if its location corresponds to a minimum of the potential and unstable if it corresponds to a maximum or a saddle point.
Rather than considering the potential, it is easier to determine the position of the light-rings by restricting to the equatorial plane and studying the stationary points of the effective potential $H(r):=-{g_{tt}}/{g_{\phi\phi}}$ that does not depend on $E$ or $L$ (see e.g., Cunha et al. (2016, 2017b); Di Filippo (2024)). This approach reveals, for example, that a single light-ring exists in the case of a Schwarzschild black hole and it is located at $r=3M$, thus marking the position of the only circular but unstable photon orbit.
Our next step consists of adding a discrete set of shells on a Schwarzschild background. In particular, in our “Gedankenexperiment” we imagine that the advanced civilisation mentioned above is able to populate the light-ring of a Schwarzschild black hole of initial mass $M_{0}$ with photons having a fine-tuned impact parameter that allows their orbits to be circular at the light-ring. We assume that the angular components of the momentum of each photon are uniformly distributed, so that the corresponding light-ring will be a 2-sphere, thus preserving the spherical symmetry of the spacetime. As the number of photons in the light-ring increases, the energy of the shell of photons will no longer be negligible, impacting the background spacetime and hence leading to a new spacetime metric that is no longer that of the Schwarzschild black hole with mass $M_{0}$. In order to obtain the resulting metric, we assume that the new metric is actually composed of a black hole and a discrete series of $N$ nested null shells of photons on (unstable) circular orbits. This allows us to consider a fully generic approach but it is trivial to restrict the relevant results to the case of a single shell, i.e., for $N=1$.
@@
where $n\in\{0,\dots,N\}$ and $\xi^{2}_{n}$ is a set of coefficients needed to guarantee the continuity between the different portions of the spacetime and where $\xi^{2}_{0}=1$. Clearly, the region $n=0$ corresponds to the section between the black hole and the first shell, while the $n$-th region corresponds to the section between the $n$-th and the $(n+1)$-th shell. Obviously, the region outside the $n$-th shell is described by a Schwarzschild spacetime with mass $M_{n}$ (note that because of our choice of time coordinate, the Killing vector $\bm{\partial_{t}}$ does not have unit norm at spatial infinity).
The mass parameter $M_{n}$ in Eq. (4) represents the total mass enclosed within the corresponding region and is thus given by $M_{n}:=M_{0}+\sum_{i=1}^{n}\delta m_{i}=M_{0}+\delta M_{n}$, where $\delta m_{i}$ indicates the mass of a single shell and $\delta M_{n}$ denotes the sum of the masses of the first $n$ shells, i.e., $\delta M_{n}:=\sum_{i=1}^{n}\delta m_{i}$.
Figure 1: Shown with a black solid line is the effective potential in the case of a discrete hierarchy of light-rings (left panel, where $M_{0}=1$, $\delta m_{1}=\delta m_{2}=0.01\,M_{0}$ and the light rings are at $r_{0}=3\,M_{0}$, $r_{1}=3\left(M_{0}+\delta m_{1}\right)$ and $r_{2}=3\left(M_{0}+\delta m_{1}+\delta m_{2}\right)$) or a continuous sequence of light-rings (right panel, where $M_{0}=1$, $R_{\rm in}=3.0\,M_{0}$ and $R_{\rm out}=3.1\,M_{0}$). The location of the light rings is shown with vertical dashed lines in the left panel and with a shaded area in the right panel. In both cases, the red dashed line shows the effective potential for a Schwarzschild black hole.
As mentioned above, $\xi^{2}_{n}$ are constants and can be fixed by simply redefining the time coordinate in each spacetime region. While this is in principle possible, it would complicate the subsequent analysis and hence we determine these constants by requiring the continuity of the time coordinate and the induced metric on the shells.
@@
whose extremal points marking the position of the unstable photon circular orbits can be easily computed after imposing $\partial_{r}H_{n}(r)=0$ and are located at the $n$ different shells with coordinate radii
$r_{n}=3M_{n}\,.$ (7)
Note that the location of the $n$-th shell coincides with the location of the light-ring of the $n-1$ spacetime region, i.e., $R_{n}=r_{n-1}$.
@@
$\xi^{2}_{1}=\frac{R_{1}-{2M_{0}}}{R_{1}-{2M_{1}}}\,,\qquad r_{1}=3M_{1}=3\left(M_{0}+\delta M_{1}\right)\,,$ (8)
@@
Is this configuration stable? In principle, to address this question we would need to perform a linear perturbation analysis taking Eq. (4) as background spacetime and assess the properties of the imaginary eigenfrequencies (see, e.g., Nagar and Rezzolla (2005)). In practice, we take a shortcut and note that it is not difficult to realise that such a spacetime is dynamically unstable. Consider, for instance, the addition of a single photon on any of the shells of the hierarchy. This will be sufficient to change the mass of the corresponding shell and hence place on unstable photon orbits all the photons on the light-rings that are at larger radii than the one on which the photon is deposited. Similarly, if a single photon is absorbed by the black hole, this would destabilise all of the light-rings, dispersing or absorbing all the photons that have been accumulated. Clearly, while we believe this conclusion is correct, a proper perturbative analysis is needed to confirm it.
@@
Continuous sequence of light-rings.  As mentioned in the introduction, the “Gedankenexperiment” can be carried out also in a different way. More specifically, rather than accumulating photons on the first light-ring till they produce a discrete contribution to the spacetime and generate a new light-ring, we can imagine a situation in which the advanced civilization concentrates on adding photons in the right amount, both in terms on numbers and impact parameters, so that the spacetime properties are varied smoothly. This would correspond to a situation in which both the mass of each individual shell and the distance among shells goes to zero while the total mass of the shells remains finite. This configuration would then describe a continuous energy distribution within a finite region of light-rings, i.e., a “light-region” with radius $r\in\left[R_{\rm in},R_{\rm out}\right]$, with $R_{\rm in}=3\,M_{0}$.
@@
Hence, we start by replacing the discrete nature of the mass in the light-ring with the continuous function $M(r)=M_{0}+\delta M(r)$, and where $\delta M(r)$ denotes the full energy of the perturbation within a radius $r$, so that $\delta M(R_{\rm in})=0$. Furthermore, as $M(r)$ denotes the total energy within a radius $r$. Also in this case we can use the coefficients $\xi_{n}$ and calculate their values in the continuous limit. More specifically, we can rewrite Eq. (5) as
$\xi^{2}_{n}=$ $\displaystyle\left(\frac{R_{n}-{2M_{n-1}}}{R_{n}-{2M_{n}}}\right)\xi^{2}_{n-1}=\prod_{i=1}^{n}\left(1+\frac{2\delta m_{i}}{R_{i}-{2M_{i}}}\right).$ (10)
The continuous limit of this product is known and it is given by a Volterra-type integral [^2] We are indebted to the anonymous referee for pointing out the possibility of employing an integral representation.. The generic expression is (see e.g. Slavík (2007))
$\lim_{\Delta X\rightarrow 0}\prod_{i=0}^{N}\left(1+f(x_{i})\Delta x\right)=\exp{\int_{x_{0}}^{x_{N}}f(x){\rm d}x}\,,$ (11)
that, in our case, yields
$\xi^{2}(r)=\exp{\displaystyle\int_{3M_{0}}^{r}\frac{2M^{\prime}(\bar{r})}{\bar{r}-2M(\bar{r})}{\rm d}\bar{r}}\,.$ ) . (12)
We can easily compute the integral using the countinuous version of Eq. (7), namely $M(r)=r/3$, obtaining
$\xi^{2}(r)=\exp{\displaystyle\int_{3M_{0}}^{r}\frac{2}{\bar{r}}{\rm d}\bar{r}}=\left(\frac{r}{3M_{0}}\right)^{2}\,.$ (13)
Therefore, the effective potential is given by
$H(r)=\begin{cases}\displaystyle\frac{1}{r^{2}}\left(1-\frac{2M_{0}}{r}\right)&r<R_{\text{in}}\\ \\ \displaystyle\frac{1}{27M_{0}^{2}}&R_{\text{in}}\leq r\leq R_{\text{out}}\\ \\ \displaystyle\frac{R_{\rm out}^{2}}{9M_{0}^{2}r^{2}}\left(1-\frac{2M\left(R_{\rm out}\right)}{r}\right)&r>R_{\text{out}}\\ \end{cases}$ (14)
@@
In summary, we have shown that it is possible to create a self-gravitating “light-region” that extends from the position of the original light-ring $R_{\rm in}$ up to $R_{\rm out}$. The corresponding potential is shown with a black solid line in the right panel of Fig. 1 and it is then easy to appreciate how we have essentially created a smooth region of light-rings where the potential is flat in the region $R_{\rm in}\leq r\leq R_{\rm out}$, hence leading to an infinite sequence of neutral circular photon orbits in such a region, i.e., with zero first and second derivative of the effective potential.
@@
We can now ask the same question: is this configuration stable? Obviously, a photon on a circular orbit well inside the light-region, i.e., at $R_{\rm in}\lesssim r\lesssim R_{\rm out}$, and that is perturbed, would still encounter in its neighbourhood other circular orbits and hence remain in the newly and perturbed location. However, it is also easy to conclude that the whole light-region is dynamically unstable after a small change in the mass of the black hole. Consider in fact that the mass of the spacetime is changed by a small but finite amount $\delta M_{0}$ without changing the mass in the light-rings, namely, consider the new mass profile $M(r)=M_{0}+\delta M_{0}+\delta M(r)$ with both $M_{0}$ and $\delta M_{0}$ constants. After straightforward computations, the new effective potential, in the range $r\in\left[R_{\rm in},R_{\rm out}\right]$, will be
$H(r)$ $=\frac{1}{27M_{0}^{2}}-\frac{2(6M_{0}-r)}{27M_{0}^{3}r}\delta M_{0}+\mathcal{O}(\delta M_{0})^{2}\,.$ (15)
So the radial derivative no longer vanishes but
$\partial_{r}H(r)=\frac{4}{9M_{0}^{2}r^{2}}\delta M_{0}+\mathcal{O}(\delta M_{0})^{2}>0\,.$ (16)
@@
Stated differently, a small perturbation in the central black hole would completely destabilise the circular orbits in the light-region, dispersing all the photons carefully accumulated there.
Finally, when looking at the left panel of Fig. 1, where each shell increases the value of the maximum of the potential, one might expect that in the continuous limit the slope of the potential would be nonzero in the shaded region of the right panel of Fig. 1. However, this is not the case, as the discrete slope $[H(r_{n})-H(r_{n-1})]/(r_{n}-r_{n-1})$ vanishes in the continuous limit. Indeed, a straightforward computation shows that $H(r_{n})-H(r_{n-1})\propto\left(\delta m_{n}\right)^{2}$ while $r_{n}-r_{n-1}\propto\delta m_{n}$, so that their ratio is zero for $\delta m_{n}\to 0$.
@@
Conclusion.  Light-rings represent a powerful tool to probe the spacetime of black holes and test different theories of gravity, both in terms of electromagnetic and gravitational radiation. So far, however, the self-gravity of light-rings has been neglected on the assumption that photons entering the light-ring are compensated by photons leaving, so that the corresponding mass is not sufficient to modify the spacetime appreciably. We have here explored the scenario in which this assumption is not true the light-ring of a Schwarzschild black hole is filled with carefully injected photons making it “self-gravitate”.